AM4 -Esercitazione 12 A.A.2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 18 dicembre 2003

Esercizio

Risolvere con la trasformata di Fourier l’equazione del calore in ℝ:

\{ \begin{matrix}u_{t} − u_{xx} = 0 x ∈ ℝ, t > 0 \\ u(0, x) = u_{0}(x) \end{matrix}

con u_{0} ∈ C^{2}, u_{0}^{(k)} ∈ R_{1}, ∀ 0 ≤ k ≤ 2.